# Quantum Pseudorandomness and Classical Complexity

### v4

@@ 5.1 $\mathsf{BQP}=\mathsf{QMA}$ relative to $(\mathcal{U},\mathcal{P})$
  • Lemma 24 and the assumption that $\mathcal{S}$ samples from a phase-invariant $\varepsilon$-approximate unitary $p(|x|)$-design imply that for any fixed $|\psi\rangle$,
    $\left|\Pr\left[V_{2}(x,|\psi\rangle)=1\right]-\Pr\left[V_{3}(x,|\psi\rangle)=1\right]\right|\leq\frac{1}{12}.$
    This is because, for each $n\in[p(|x|)]\setminus[d]$ and $m\in\{0,1\}^{n}$, switching $U_{nm}$ between the Haar measure and the approximate design changes the acceptance probability by at most $\varepsilon=\frac{1}{12p(|x|)2^{p(|x|)}}$. Summing over all these changes, the total difference in acceptance probability is at most
    $\sum_{n=d+1}^{p(|x|)}\sum_{m\in\{0,1\}^{n}}\frac{1}{12p(|x|)2^{p(|x|)}}\leq\frac{1}{12}.$
    This in turn implies that $\left|\mathrm{acc}(V_{2})-\mathrm{acc}(V_{3})\right|\leq\frac{1}{12}$.
  • Zhandry [Zha12, Theorem 6.1] shows that a quantum algorithm that makes $T$ queries to a random function can be exactly simulated by the same algorithm with $T$ queries to a $2T$-wise independent function, so $\mathrm{acc}(V_{3})=\mathrm{acc}(V_{4})$.
  • Because $||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}\leq\frac{1}{12p(|x|)}$ for each $n\in[d]$ with probability at least $\frac{2}{3}$ over $\mathcal{A}$, from 6 and because $\mathcal{V}$ makes at most $p(|x|)$ queries, it holds that $\Pr_{\mathcal{A}}\left[|\mathrm{acc}(V_{4})-\mathrm{acc}(V_{5})|\geq\frac{1}{12}\right]\leq\frac{1}{3}$.
Putting these bounds together, we have that, except with probability $2e^{-|x|}$ over $\mathcal{U}$:
@@
$\Pi(x)=1\quad$ $\implies\quad\Pr_{\mathcal{A}}\left[\max_{|\psi\rangle}\Pr\left[\widetilde{\mathcal{V}}(x,|\psi\rangle)=1\right]\geq\frac{2}{3}\right]\geq\frac{2}{3}$
@@
$\Pi(x)=0\quad$ $\implies\quad\Pr_{\mathcal{A}}\left[\max_{|\psi\rangle}\Pr\left[\widetilde{\mathcal{V}}(x,|\psi\rangle)=1\right]\leq\frac{1}{3}\right]\geq\frac{2}{3}.$
@@
This is to say that $\mathcal{A}$ correctly decides $\Pi(x)$, expect with probability at most $2e^{-|x|}$ over $\mathcal{U}$. By the Borel–Cantelli lemma, because $\sum_{i=1}^{\infty}2^{i}\cdot 2e^{-i}=\frac{4}{e-2}<\infty$, $\mathcal{A}$ correctly decides $\Pi(x)$ for all but finitely many $x\in\mathrm{Dom}(\Pi)$, with probability $1$ over $\mathcal{U}$. Hence, with probability $1$ over $\mathcal{U}$, $\mathcal{A}$ can be modified into an algorithm $\mathcal{A}^{\prime}$ that agrees with $\Pi$ on every $x\in\mathrm{Dom}(\Pi)$, by simply hard-coding those $x$ on which $\mathcal{A}$ and $\Pi$ disagree.
@@
Because there are only countably many $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$ machines, we can union bound over all $\Pi\in\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$ to conclude that $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}\subseteq\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}$ with probability $1$ over $\mathcal{U}$. ∎
5.2 Pseudorandom unitaries relative to $(\mathcal{U},\mathcal{P})$
@@
We proceed to the second part of the oracle construction, showing that PRUs exist relative to $(\mathcal{U},\mathcal{P})$. We begin with a lemma establishing that the average advantage of a polynomial-time adversary is small against our PRU construction. Here, we should think of $\{U_{k}\}_{k\in[N]}$ as the PRU ensemble.
@@
Consider a quantum algorithm $\mathcal{A}^{O,U}$ that makes $T$ queries to $O\in\mathbb{U}(D)$ and $U=(U_{1},\ldots,U_{N})\in\mathbb{U}(D)^{N}$. For fixed $U$, define:
@@
The next lemma uses Lemma 30 to show that the advantage of $\mathcal{A}$ is small with extremely high probability, which follows from the strong concentration properties of the Haar measure (Theorem 9).
@@
Consider a quantum algorithm $\mathcal{A}^{O,U}$ that makes $T$ queries to $O\in\mathbb{U}(D)$ and $U=(U_{1},\ldots,U_{N})\in\mathbb{U}(D)^{N}$. Let $\mathrm{adv}(\mathcal{A}^{U})$ be defined as in Lemma 31. Then there exists a universal constant $c>0$ such that for any $p$,
$\Pr_{U\sim\mu_{D}^{N}}\left[\left|\mathrm{adv}(\mathcal{A}^{U})\right|\geq p\right]\leq 2\exp\left(-\frac{(D-2)\left(p-cT^{2}/N\right)^{2}}{384T^{2}}\right).$
@@
By Lemma 27, $\mathrm{adv}(\mathcal{A}^{U})$ is $4T$-Lipschitz as a function of $U$, because $\mathrm{adv}(\mathcal{A}^{U})$ can be expressed as the the difference between the acceptance probabilities of two algorithms that each make $T$ queries to $U$. Combining Lemma 30 and Theorem 9, we obtain:
$\Pr_{U\sim\mu_{D}^{N}}\left[\mathrm{adv}(\mathcal{A}^{U})\geq p\right]\leq\exp\left(-\frac{(D-2)\left(p-cT^{2}/N\right)^{2}}{384T^{2}}\right).$
@@
Completing the security proof of the pseudorandom unitary construction amounts to combining Lemma 31 with a union bound over all possible polynomial-time adversaries.
@@
Let $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$ be a quantum oracle where each $\mathcal{U}_{n}$ is chosen randomly from $\mu_{2^{n}}^{2^{n}}$. Let $\mathcal{P}$ be an arbitrary $\mathsf{PSPACE}$-complete language. Then with probability $1$ over $\mathcal{U}$, there exists a family of PRUs relative to $(\mathcal{U},\mathcal{P})$ with $n(\kappa)=\kappa$.
@@
Fix an input length $n\in\mathbb{N}$. We take the key set $\{0,1\}^{\kappa}=[2^{n}]$ and take the PRU family to be $\{U_{k}\}_{k\in\{0,1\}^{\kappa}}$, where $\mathcal{U}_{n}=(U_{1},U_{2},\ldots,U_{2^{n}})\in\mathbb{U}(2^{n})^{2^{n}}$. In words, the family consists of the $2^{n}$ different Haar-random $n$-qubit unitaries specified by $\mathcal{U}_{n}$.
@@
Without loss of generality, assume the adversary is a uniform polynomial-time quantum algorithm $\mathcal{A}^{O,\mathcal{U},\mathcal{P}}(1^{n},x)$, where $x\in\{0,1\}^{\mathrm{poly}(n)}$ is the advice and $O\in\mathbb{U}(2^{n})$ is the oracle that the adversary seeks to distinguish as pseudorandom or Haar-random.
@@
By Lemma 31 with $N=D=2^{n}$ and $T=\mathrm{poly}(n)$, for any fixed $x\in\{0,1\}^{\mathrm{poly}(n)}$, $\mathcal{A}^{O,\mathcal{U},\mathcal{P}}(1^{n},x)$ achieves non-negligible advantage with extremely low probability over $\mathcal{U}$. This is to say that for any $p=\frac{1}{\mathrm{poly}(n)}$:
$\Pr_{\mathcal{U}_{n}\sim\mu_{2^{n}}^{2^{n}}}\left[\left|\Pr_{k\sim[2^{n}]}\left[\mathcal{A}^{U_{k},\mathcal{U},\mathcal{P}}(1^{n},x)=1\right]-\Pr_{O\sim\mu_{2^{n}}}\left[\mathcal{A}^{O,\mathcal{U},\mathcal{P}}(1^{n},x)=1\right]\right|\geq p\right]\leq\exp\left(-\frac{2^{n}}{\mathrm{poly}(n)}\right).$
@@
By a union bound over all $x\in\{0,1\}^{\mathrm{poly}(n)}$, $\mathcal{A}^{O,\mathcal{U},\mathcal{P}}(1^{n},x)$ achieves advantage larger than $p$ for any $x\in\{0,1\}^{\mathrm{poly}(n)}$ with probability at most $2^{\mathrm{poly}(n)}\cdot\exp\left(-\frac{2^{n}}{\mathrm{poly}(n)}\right)\leq\mathrm{negl}(n)$. Hence, by the Borel–Cantelli lemma, $\mathcal{A}$ achieves negligible advantage for all but finitely many input lengths $n\in\mathbb{N}$ with probability $1$ over $\mathcal{U}$, as $\sum_{n=1}^{\infty}\mathrm{negl}(n)<\infty$. This is to say that $\{U_{k}\}_{k\in\{0,1\}^{\kappa}}$ defines a PRU ensemble. ∎
@@
We expect that using the techniques of Chung, Guo, Liu, and Qian [CGLQ20], one can extend Theorem 32 to a security proof against adversaries with quantum advice. Some version of [CGLQ20, Theorem 5.14] likely suffices. The idea is that breaking the PRU should remain hard even if $\mathcal{A}$ could query an explicit description of $O$ and explicit descriptions of $U_{k}$ for $k\in[2^{n}]$, which is a strictly more powerful model. But then this corresponds to the security game defined in [CGLQ20, Definition 5.12], except that the range of the random oracle is $\mathbb{U}(D)$ rather than the finite set $[M]$.
Acknowledgements

### v5

@@ 5.2 $\mathsf{PromiseBQP}=\mathsf{PromiseQMA}$ relative to $(\mathcal{U},\mathcal{C})$
Because $\mathcal{W}$ calls $\widetilde{\mathcal{U}}$ at most $p(|x|)$ times, and because diamond distance between unitary channels is preserved under taking inverses, 7 implies that for any $|\psi\rangle$,
@@
${\left|\Pr\left[\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\widetilde{\mathcal{U}};|\psi\rangle)=1\right]-\Pr\left[\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\mathcal{U};|\psi\rangle)=1\right]\right|}\leq\frac{p(|x|)}{2}||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}.$
Hence, we also have
$\left|f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})-f(\mathcal{U},\overline{\mathcal{U}})\right|$ $={\left|\max_{|\psi\rangle}\Pr\left[\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\widetilde{\mathcal{U}};|\psi\rangle)=1\right]-\max_{|\psi\rangle}\Pr\left[\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\mathcal{U};|\psi\rangle)=1\right]\right|}$
$\leq\frac{p(|x|)}{2}||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond},$
and therefore, by Jensen’s inequality,
${\left|\mathop{\mathbb{E}}_{\overline{U}\sim\mathcal{D}}[f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})]-\mathop{\mathbb{E}}_{\overline{U}\sim\mathcal{D}}[f(\mathcal{U},\overline{\mathcal{U}})]\right|}\leq\frac{p(|x|)}{2}||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}.$
Because the estimates $\widetilde{\mathcal{U}_{n}}$ satisfy $||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}\leq\frac{1}{6p(|x|)}$ with probability at least $\frac{2}{3}$ over the randomness of $\mathcal{A}$, we see:
@@
$\Pr_{\mathcal{A}}\left[\left|\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}\left[f(\mathcal{U},\overline{\mathcal{U}})\right]-\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}\left[f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})\right]\right|\geq\frac{1}{12}\right]\leq\frac{1}{3}.$
Combining with Equation 4, and recalling the acceptance criterion of $\mathcal{A}$ from Equation 2, we conclude that except with probability at most $2e^{-|x|}$ over $\mathcal{U}$,
$\Pr_{\mathcal{A}}\left[\left|f(\mathcal{U},\mathcal{U})-\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}\left[f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})\right]\right|\geq\frac{1}{6}\right]\leq\frac{1}{3}.$
So, except with probability $2e^{-|x|}$ over $\mathcal{U}$:
$\Pi(x)=1\quad$ $\implies\quad f(\mathcal{U},\mathcal{U})\geq\frac{11}{12}\quad\implies\quad\Pr\left[\mathcal{A}^{\mathcal{U},\mathcal{C}}(x)=1\right]\geq\frac{2}{3}$
$\Pi(x)=0\quad$ $\implies\quad f(\mathcal{U},\mathcal{U})\leq\frac{1}{12}\quad\implies\quad\Pr\left[\mathcal{A}^{\mathcal{U},\mathcal{C}}(x)=0\right]\geq\frac{2}{3}.$
@@
This is to say that $\mathcal{A}$ correctly decides $\Pi(x)$, expect with probability at most $2e^{-|x|}$ over $\mathcal{U}$. By the Borel–Cantelli lemma (Lemma 6), because $\sum_{i=1}^{\infty}2^{i}\cdot 2e^{-i}=\frac{4}{e-2}<\infty$, $\mathcal{A}$ correctly decides $\Pi(x)$ for all but finitely many $x\in\mathrm{Dom}(\Pi)$, with probability $1$ over $\mathcal{U}$. Hence, with probability $1$ over $\mathcal{U}$, $\mathcal{A}$ can be modified into an algorithm $\mathcal{A}^{\prime}$ that agrees with $\Pi$ on every $x\in\mathrm{Dom}(\Pi)$, by simply hard-coding those $x$ on which $\mathcal{A}$ and $\Pi$ disagree.
@@
Because there are only countably many $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$ machines, we can union bound over all $\Pi\in\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$ to conclude that $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}\subseteq\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{C}}$ with probability $1$. ∎
5.3 Pseudorandom unitaries relative to $(\mathcal{U},\mathcal{C})$
We proceed to the second part of the oracle construction, showing that PRUs exist relative to $(\mathcal{U},\mathcal{C})$. In fact, the security proof will not depend on $\mathcal{C}$: the same PRU construction is secure for any language $\mathcal{C}$ that is independent of the randomly sampled $\mathcal{U}$. The PRU ensemble for a given length is supplied directly by $\mathcal{U}$. That is, for a given length $n$, the PRU ensemble is uniform over the $2^{n}$ different $n$-qubit unitaries in $\mathcal{U}_{n}$.
@@ 5.3 Pseudorandom unitaries relative to $(\mathcal{U},\mathcal{C})$
We begin with a lemma establishing that the average advantage of a polynomial-time adversary is small against our PRU construction. Here, we should think of $\{U_{k}\}_{k\in[N]}$ as the PRU ensemble.
@@
Consider a quantum algorithm $\mathcal{A}^{O,U}$ that makes $T$ queries to $U=(U_{1},\ldots,U_{N})\in\mathbb{U}(D)^{N}$ and $O\in\mathbb{U}(D)$. For fixed $U$, define:
@@
The next lemma uses Lemma 31 to show that the advantage of $\mathcal{A}$ is small with extremely high probability, which follows from the strong concentration properties of the Haar measure (Theorem 10). This strengthening of Lemma 31 will be needed to argue that the advantage remains small even after union bounding over all choices of the classical advice.
@@
Consider a quantum algorithm $\mathcal{A}^{O,U}$ that makes $T$ queries to $U=(U_{1},\ldots,U_{N})\in\mathbb{U}(D)^{N}$ and $O\in\mathbb{U}(D)$. Let $\mathrm{adv}(\mathcal{A}^{U})$ be defined as in Lemma 31. Then there exists a universal constant $c>0$ such that for any $p\geq cT^{2}/N$,
$\Pr_{U\sim\mu_{D}^{N}}\left[\left|\mathrm{adv}(\mathcal{A}^{U})\right|\geq p\right]\leq 2\exp\left(-\frac{(D-2)\left(p-cT^{2}/N\right)^{2}}{96T^{2}}\right).$
@@
By Lemma 28, $\mathrm{adv}(\mathcal{A}^{U})$ is $2T$-Lipschitz as a function of $U$, because $\mathrm{adv}(\mathcal{A}^{U})$ can be expressed as the difference between the acceptance probabilities of two algorithms that each make $T$ queries to $U$. Combining Lemma 31 and Theorem 10, we obtain:
$\Pr_{U\sim\mu_{D}^{N}}\left[\mathrm{adv}(\mathcal{A}^{U})\geq p\right]\leq\exp\left(-\frac{(D-2)\left(p-cT^{2}/N\right)^{2}}{96T^{2}}\right).$
@@
Completing the security proof of the pseudorandom unitary construction amounts to combining Lemma 32 with the aforementioned union bound over all possible polynomial-time adversaries.
@@
Let $\mathcal{C}$ be any fixed language. Then with probability $1$ over $\mathcal{U}\sim\mathcal{D}$, there exists a family of PRUs relative to $(\mathcal{U},\mathcal{C})$ with $n(\kappa)=\kappa$.
@@
Fix an input length $n\in\mathbb{N}$. We take the key set $\{0,1\}^{\kappa}=\{0,1\}^{n}\equiv[2^{n}]$ and take the PRU family to be $\{U_{k}\}_{k\in\{0,1\}^{n}}$, where $\mathcal{U}_{n}=(U_{1},U_{2},\ldots,U_{2^{n}})\in\mathbb{U}(2^{n})^{2^{n}}$. In words, the family consists of the $2^{n}$ different Haar-random $n$-qubit unitaries supplied by $\mathcal{U}_{n}$. Note that this family of unitaries has an efficient implementation relative to the oracle. This is because we can simulate an application of $U_{k}$ to some $n$-qubit $|\psi\rangle$ using one query to $\mathcal{U}_{n}$, via $\mathcal{U}_{n}|k\rangle|\psi\rangle=(I\otimes U_{k})|k\rangle|\psi\rangle$. So, it remains only to show the computational indistinguishability criterion of Definition 16.
@@
Without loss of generality, assume the adversary is a uniform polynomial-time quantum algorithm $\mathcal{A}^{O,\mathcal{U},\mathcal{C}}(1^{n},x)$, where $x\in\{0,1\}^{\mathrm{poly}(n)}$ is the advice and $O\in\mathbb{U}(2^{n})$ is the oracle that the adversary seeks to distinguish as pseudorandom or Haar-random.
@@
By Lemma 32 with $N=D=2^{n}$ and $T=\mathrm{poly}(n)$, for any fixed $x\in\{0,1\}^{\mathrm{poly}(n)}$, $\mathcal{A}^{O,\mathcal{U},\mathcal{C}}(1^{n},x)$ achieves non-negligible advantage with extremely low probability over $\mathcal{U}$. (The additional oracle $\mathcal{C}$ has no effect on the query complexity result because it fixed and independent of $\mathcal{U}$.) This is to say that for any $p=\frac{1}{\mathrm{poly}(n)}$:
$\Pr_{\mathcal{U}_{n}\sim\mu_{2^{n}}^{2^{n}}}\left[\left|\Pr_{k\sim[2^{n}]}\left[\mathcal{A}^{U_{k},\mathcal{U},\mathcal{C}}(1^{n},x)=1\right]-\Pr_{O\sim\mu_{2^{n}}}\left[\mathcal{A}^{O,\mathcal{U},\mathcal{C}}(1^{n},x)=1\right]\right|\geq p\right]\leq\exp\left(-\frac{2^{n}}{\mathrm{poly}(n)}\right).$
@@
By a union bound over all $x\in\{0,1\}^{\mathrm{poly}(n)}$, $\mathcal{A}^{O,\mathcal{U},\mathcal{C}}(1^{n},x)$ achieves advantage larger than $p$ for any $x\in\{0,1\}^{\mathrm{poly}(n)}$ with probability at most $2^{\mathrm{poly}(n)}\cdot\exp\left(-\frac{2^{n}}{\mathrm{poly}(n)}\right)\leq\mathrm{negl}(n)$. Hence, by the Borel–Cantelli lemma (Lemma 6), $\mathcal{A}$ achieves negligible advantage for all but finitely many input lengths $n\in\mathbb{N}$ with probability $1$ over $\mathcal{U}$, as $\sum_{n=1}^{\infty}\mathrm{negl}(n)<\infty$. This is to say that $\{U_{k}\}_{k\in\{0,1\}^{n}}$ defines a PRU ensemble. ∎
@@
We expect that using the techniques of Chung, Guo, Liu, and Qian [CGLQ20], one can extend Theorem 33 to a security proof against adversaries with quantum advice. Some version of [CGLQ20, Theorem 5.14] likely suffices. The idea is that breaking the PRU should remain hard even if $\mathcal{A}$ could query an explicit description of $O$ and explicit descriptions of $U_{k}$ for $k\in[2^{n}]$, which is a strictly more powerful model. But then this corresponds to the security game defined in [CGLQ20, Definition 5.12], except that the range of the random oracle is $\mathbb{U}(D)$ rather than the finite set $[M]$. Perhaps a sufficiently fine discretization of $\mathbb{U}(D)$ would suffice to apply the [CGLQ20] framework. We believe this is doable but tedious, and leave it to future work.
5.4 Alternative oracles
An earlier version of this paper [Kre21a] claimed to show the same results, Theorems 30 and 33, but relative to a different oracle. Instead of $\mathcal{O}=(\mathcal{U},\mathcal{C})$ for $\mathcal{U}\sim\mathcal{D}$, the earlier oracle used a different classical language in place of $\mathcal{C}$; the oracle chosen was $\mathcal{O}=(\mathcal{U},\mathcal{P})$ where $\mathcal{P}$ is an arbitrary $\mathsf{PSPACE}$-complete language. As noted above, PRUs still exist relative to this oracle because Theorem 33 works regardless of the choice of classical language. However, the claim that $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$ contained a bug in the proof. This incorrect step amounted to conflating the two quantities
$\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\max_{|\psi\rangle}\Pr\left[\mathcal{V}^{\mathcal{U},\mathcal{P}}(|\psi\rangle)=1\right]\right]$ (5)
and
$\max_{|\psi\rangle}\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\Pr\left[\mathcal{V}^{\mathcal{U},\mathcal{P}}(|\psi\rangle)=1\right]\right],$
which are not the same. Nevertheless, we conjecture that the previous proof can be restored:
Conjecture 34.
With probability $1$ over $\mathcal{U}\sim\mathcal{D}$, $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$, where $\mathsf{P}$ is an arbitrary $\mathsf{PSPACE}$-complete language.
A careful inspection of [Kre21a] reveals that 34 could be proved by showing that the quantity in Equation 5 is approximable in $\mathsf{PSPACE}$. We see a possible approach to establishing this, which relies on the following well-known analogue of the polynomial method [BBC+01] for $\mathsf{QMA}$ verifiers:
Proposition 35 (Proved in [Aar09, Lemma 4]).
Let $\mathcal{V}$ be a $\mathsf{QMA}$-verifier that receives an $m$-qubit witness and makes $T$ queries to a unitary $\mathcal{O}$. Then there exists a matrix-valued polynomial $M(\mathcal{O})$ of degree $2T$ in $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ such that for any $m$-qubit $|\psi\rangle$,
$\langle\psi|M(\mathcal{O})|\psi\rangle=\Pr\left[\mathcal{V}^{\mathcal{O}}(|\psi\rangle)=1\right].$
Proof.
Without loss of generality, suppose that on input $|\psi\rangle$, $\mathcal{V}$ appends $n$ ancilla qubits initialized to $|0\rangle$, applies a unitary $U(\mathcal{O})$ that may involve queries to $\mathcal{O}$, and then measures the first qubit. Then the matrix $M(\mathcal{O})$ is:
$(I\otimes\langle 0^{n}|)U(\mathcal{O})^{\dagger}(|1\rangle\langle 1|\otimes I)U(\mathcal{O})(I\otimes|0^{n}\rangle),$
which clearly satisfies
$\langle\psi|M(\mathcal{O})|\psi\rangle=\Pr\left[\mathcal{V}^{\mathcal{O}}(|\psi\rangle)=1\right].$
Additionally, the entries of $M(\mathcal{O})$ are polynomials of degree $2T$ in $\mathcal{O}$ and $\mathcal{O}^{\dagger}$ because $U(\mathcal{O})$ is a polynomial of degree $T$ [BBC+01], and $U(\mathcal{O})$ appears twice in the expression. ∎
A key observation is the following: if $p(\mathcal{O})\coloneqq\max_{|\psi\rangle}\Pr\left[\mathcal{V}^{\mathcal{O}}(|\psi\rangle)=1\right]$, then for any $k\in\mathbb{N}$
$p(\mathcal{O})^{k}\leq\mathrm{Tr}(M(\mathcal{O})^{k})\leq 2^{m}p(\mathcal{O})^{k}.$
Equivalently,
$\frac{\mathrm{Tr}(M(\mathcal{O})^{k})^{1/k}}{2^{m/k}}\leq p(\mathcal{O})\leq\mathrm{Tr}(M(\mathcal{O})^{k})^{1/k}.$
So, by choosing $k$ to be sufficiently large (say, $100m$), $\mathrm{Tr}(M(\mathcal{O})^{k})^{1/k}$ provides an arbitrarily precise estimate of $p(\mathcal{O})$. Thus, to approximate Equation 5, it suffices to approximate
$\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\mathrm{Tr}(M(\mathcal{U},\mathcal{P})^{k})^{1/k}\right],$
which we believe is achievable in $\mathsf{PSPACE}$. We first observe that, as a consequence of the concentration of the Haar measure (Theorem 10), the above quantity should satisfy
$\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\mathrm{Tr}(M(\mathcal{U},\mathcal{P})^{k})^{1/k}\right]\approx\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\mathrm{Tr}(M(\mathcal{U},\mathcal{P})^{k})\right]^{1/k},$
as long as $\mathcal{V}$ only makes queries to $\mathcal{U}$ in sufficiently large dimension.
Notice that $\mathrm{Tr}(M(\mathcal{U},\mathcal{P})^{k})$ is a polynomial of degree $2Tk$ in the entries of $\mathcal{U}$ and $\mathcal{P}$ (and their inverses). Moreover, the proof of Proposition 35 reveals that the coefficients of this polynomial are computable in $\mathsf{PSPACE}$, by standard path integral techniques [NC10, Section 4.5.5]. The main question, then, is whether one can average this polynomial over the Haar measure in $\mathsf{PSPACE}$. With some additional work, we believe this could be established via either
  1. (1)
    Showing that the Weingarten calculus [CMN22], used for evaluating Haar integrals, is computable in $\mathsf{PSPACE}$, or
  2. (2)
    Proving that a sufficiently strong notion of unitary $t$-design (Definition 18) yields an approximation of $\mathop{\mathbb{E}}_{\mathcal{U}\sim\mathcal{D}}\left[\mathrm{Tr}(M(\mathcal{U},\mathcal{P})^{k})\right]$. The challenge here is that this expression involves applications of both $\mathcal{U}$ and $\mathcal{U}^{\dagger}$, even when $\mathcal{V}$ only makes queries to $\mathcal{U}$ in the forward direction. So, Lemma 25 does not seem applicable.
Acknowledgments